the width of a rectangle is two meters more then one third of the length. The perimeter of the rectangle is forty four meters. Find the measure of each angle.

Answer by mananth(16946)

the width of a rectangle is two meters more then one third of the length. The perimeter of the rectangle is forty four meters. Find the measure of each angle.
...
let length = l
l/3 +2 = width
...
2*(l+w)=perimeter
2*(l+l/3+2)=40
/2
l+l/3 +2 = 20
l+l/3 = 20-2
l+l/3 = 18
4l/3=18
multiply by 3
4l = 18*3
41=54
/4
l = 54/4
l=13.5 meters ---- The length
width = (1/3*13.5) +2
width = 6.5 meters
...
CHECK
2*(13.5+6.5)=
2*(20)=40 meters
...
m.ananth@hotmail.ca